The common stock of the company sells for $33 a share, has a rate of return of 11.140%, and a dividend growth rate of 2% annually. what was the amount of the last dividend paid? please tell me in detail how to solve and show the formula used.
Answer #1

P0 = D0(1 + g) / (r - g)

$33 = D0(1.02) / (0.11140 - 0.02)

$33 = D0(1.02) / 0.0914

$33 × 0.0914 = D0(1.02)

D0(1.02) = $3.0162

D0 = $3.0162 / 1.02

D0 = $2.96